# Some frameworks for dissipative evolution in multiscale non-equilibrium thermodynamics

### v1

@@ Abstract
In this paper, we review and compare some frameworks for dissipation in non-equilibrium thermodynamics. We start with a brief overview of classical irreversible thermodynamics and gradient dynamics. Then we discuss several specific frameworks including Rayleigh dissipation potential and the dissipative d’Alembert framework, showing their relations with gradient dynamics. Finally, we discuss frameworks for dissipative evolution generated from Poisson brackets.
@@ 1 Introduction
Historically, the first passage from the microscopic particle dynamics to classical equilibrium thermodynamics was made by Ludwig Boltzmann [23] for an ideal gas. The Boltzmann entropy introduced in the passage indeed plays the role of a Lyapunov function for the approach to thermodynamic equilibrium states, and if the entropy is evaluated at the approached equilibrium states, it becomes the fundamental equilibrium thermodynamic relation for the ideal gas. In the Gibbs passage [24] from the microscopic particle dynamics to classical equilibrium thermodynamics (which applies to all macroscopic systems), the time evolution, in the course of which the Gibbs entropy is maximized, is absent. Many frameworks have been developed for the passage from fluid mechanics to classical equilibrium thermodynamics. Fluid mechanics (classical and extended) is supplemented with the assumption of local equilibrium (classical and extended) and the requirement that the entropy plays the role of a Lyapunov function in the passage (second law of thermodynamics). Among the frameworks, we mention Classical Irreversible Thermodynamics (CIT) [57, 58, 13, 65], Extended Irreversible Thermodynamics [43],[48], Internal Variables Thermodynamics [5], Rational Thermodynamics [71], Rational Extended Thermodynamics [55].
@@
In this paper, we focus on geometric frameworks for multiscale dissipative time evolution. Such frameworks are formulated in the multiscale setting (i.e., they can be used in the context of fluid mechanics, kinetic theory, and other autonomous mesoscopic theories) and provide a unification and additional physical and mathematical insights. Among the geometric frameworks, we discuss generalized gradient dynamics [62, 39], the General Equation for Non-Equilibrium Reversible-Irreversible Coupling (GENERIC)[31],[30], [38, 60, 59, 63], [6], and metriplectic systems [53, 44, 31], where gradient dynamics is combined with Hamiltonian mechanics. In the case when the target mesoscopic theory is not equilibrium thermodynamics but a mesoscopic theory involving fewer details, the role of the entropy is replaced by the Rayleigh dissipation potential [67, 47, 15]. Dissipative terms can be also added as constraints in the variational formulation of mechanics, leading to the d’Alembert variational principle in non-equilibrium thermodynamics [21]. Finally, when we have a mechanical (Hamiltonian) system, dissipative evolution can be constructed from the underlying Poisson bracket and from the Hamiltonian (double bracket dissipation [8] and Ehrenfest regularization [64]). In particular, we show that the Rayleigh-potential framework can be cast into the GENERIC form, employing generalized gradient dynamics as the dissipative part of the evolution. We also formulate a version of the dissipative d’Alembert variational principle using the Rayleigh dissipation potential, which makes it easier to eliminate fast-relaxing variables in a geometric manner. Finally, we show a variational formulation of the GENERIC framework.
@@
In Section 2, we motivate the frameworks for dissipative evolution by pointing out some problems of CIT. Section 3 contains generalized gradient dynamics while Section 4 is about the Rayleigh dissipation potential. In Section 5 we discuss the dissipative d’Alembert variational principle, and finally in Section 6 we recall some framework where dissipative evolution is constructed from the Poisson bracket the considered system.
@@ 2 Classical irreversible thermodynamics
State variables in the context of Classical Irreversible Thermodynamics [13] are chosen to be $\mathbf{x}=(\mathbf{q},e)$, where $e$ is the volumetric total energy density and $\mathbf{q}$ represents the remaining variables (for instance mass density or momentum density). The time evolution of the volumetric energy density $e$ is governed by
@@
where $\mathbf{J}_{e}$ is the energy flux (or heat flux). This evolution equation expresses the conservation of the total energy. The evolution of the remaining state variables $\mathbf{q}$ is given by a force-flux constitutive relation based on the entropy production formula,
@@
$\partial_{t}s(\mathbf{x})=$ $\frac{\partial s}{\partial\mathbf{x}}\cdot\partial_{t}\mathbf{x}=\frac{1}{T}\left(-\nabla\cdot\mathbf{J}_{e}\right)+\frac{\partial s}{\partial\mathbf{q}}\cdot\partial_{t}\mathbf{q}$
@@
$=$ $-\nabla\cdot\left(\frac{1}{T}\mathbf{J}_{e}\right)\underbrace{+\mathbf{J}_{e}\cdot\nabla\frac{1}{T}+\partial_{t}\mathbf{q}\cdot\frac{\partial s}{\partial\mathbf{q}}}_{\geq 0},$ (2)
where $s$ is the volumetric entropy density, $T=\left(\frac{\partial s}{\partial e}\right)^{-1}$ is the temperature.
In the particular case when we expect the remaining state variables $\mathbf{q}$ to be conserved, that is
$\partial_{t}\mathbf{q}=-\nabla\cdot\mathbf{J}_{\mathbf{q}},$ (3)
the balance of entropy becomes
@@
$\partial_{t}s=-\nabla\cdot\left(\frac{1}{T}\mathbf{J}_{e}+\mathbf{J}_{\mathbf{q}}\frac{\partial s}{\partial\mathbf{q}}\right)+\underbrace{\mathbf{J}_{e}\cdot\nabla\frac{1}{T}+\mathbf{J}_{\mathbf{q}}\cdot\nabla\frac{\partial s}{\partial\mathbf{q}}}_{\geq 0}.$ (4)
@@
To guarantee the non-negativity of the entropy production, the fluxes are then expressed as linear combinations of the thermodynamic forces (gradients of the derivatives of entropy),
@@
$\begin{pmatrix}\mathbf{J}_{e}\\ \mathbf{J}_{\mathbf{q}}\end{pmatrix}=\mathbf{M}(\mathbf{x})\cdot\begin{pmatrix}\nabla\frac{\partial s}{\partial e}\\ \nabla\frac{\partial s}{\partial\mathbf{q}}\end{pmatrix}$ (5)
@@
where $\mathbf{M}$ is a linear operator that is positive semi-definite, $\mathbf{x}\cdot\mathbf{M}(\mathbf{x})\cdot\mathbf{x}\geq 0$.
If a state variable is not conserved, the evolution equation for entropy becomes
$\partial_{t}s=-\nabla\cdot\left(\frac{1}{T}\mathbf{J}_{e}\right)+\underbrace{\mathbf{J}_{e}\cdot\nabla\frac{1}{T}+\partial_{t}\mathbf{q}\cdot\frac{\partial s}{\partial\mathbf{q}}}_{\geq 0},$ (6)
which leads to the closure
$\begin{pmatrix}\mathbf{J}_{e}\\ \partial_{t}\mathbf{q}\end{pmatrix}=\mathbf{M}(\mathbf{x})\cdot\begin{pmatrix}\nabla\frac{\partial s}{\partial e}\\ \frac{\partial s}{\partial\mathbf{q}}\end{pmatrix}$. (7)
@@
It then depends on the particular problem how the non-negativity of the entropy production is guaranteed. Often, the entropy is chosen as quadratic in the non-conserved state variables (for instance heat flux in the Cattaneo equation [12]), which leads to Extended Irreversible Thermodynamics [43] and Internal Variable Thermodynamics [5, 4, 51].
@@
  • Evolution of state variables $\mathbf{q}$ does not in general have to be in the divergence form, as in the case of polymeric fluids (conformation tensor is not conserved) [7], magnetohydrodynamics (where the Godunov-Powell terms are important for numerical stability [26]), Reynolds stress in turbulence [37], and others [63]. Instead, the evolution equation for $\mathbf{q}$ can have a part that is in the divergence form and a part that is not. Moreover, we can move some terms between the two parts, which makes the identification of fluxes and forces ambiguous.
@@ 3.1 Generalized gradient dynamics
where $\Xi$ is a dissipation potential, $S$ is the entropy, and $\mathbf{x}^{*}$ are thermodynamic forces conjugate to $\mathbf{x}$. The derivative $\delta\bullet/\delta\bullet$ stands for the functional derivative so that we can differentiate functionals with respect to fields. The dissipation potential is a non-negative functional of $\mathbf{x}^{*}$ that has a minimum at $\mathbf{x}^{*}=0$, is convex in the vicinity of the minimum, and is radially monotonous, $\mathbf{x}^{*}\cdot\frac{\delta\Xi}{\delta\mathbf{x}^{*}}$; see [68, 42]. The dissipation potential $\Xi(\mathbf{x},\mathbf{x}^{*})$ also depends on the state variables $\mathbf{x}$.

### v2

# Comparison of some geometric frameworks for dissipative evolution in multiscale non-equilibrium thermodynamics

@@ Abstract
In this paper, we review and compare some geometric frameworks for dissipation in non-equilibrium thermodynamics. We start with a brief overview of classical irreversible thermodynamics and gradient dynamics. Then we discuss several specific frameworks including Rayleigh dissipation potential and the dissipative d’Alembert framework, showing their relations with gradient dynamics. Finally, we discuss frameworks for dissipative evolution generated from Poisson brackets.
@@ 1 Introduction
Historically, the first passage from the microscopic particle dynamics to classical equilibrium thermodynamics was made by Ludwig Boltzmann [34] for an ideal gas. The Boltzmann entropy introduced in the passage indeed plays the role of a Lyapunov function for the approach to thermodynamic equilibrium states, and if the entropy is evaluated at the approached equilibrium states, it becomes the fundamental equilibrium thermodynamic relation for the ideal gas. In the Gibbs passage [35] from the microscopic particle dynamics to classical equilibrium thermodynamics (which applies to all macroscopic systems), the time evolution, in the course of which the Gibbs entropy is maximized, is absent. Many frameworks have been developed for the passage from fluid mechanics to classical equilibrium thermodynamics. Fluid mechanics (classical and extended) is supplemented with the assumption of local equilibrium (classical and extended) and the requirement that the entropy plays the role of a Lyapunov function in the passage (second law of thermodynamics).
Technically speaking, if we do not restrict the stability analysis to the reachable submanifold of the whole space of state variables (states with the same energy, momentum, etc., as the equilibrium state), the entropy is rather an S-function [4] than a Lyapunov function [63]. However, for simplicity, we shall call it a Lyapunov function, tacitly assuming the restriction to the reachable submanifold, which is often reflected in the choice of vector spaces in the case of infinite-dimensional systems [14].
Another technical issue that we are not addressing in this paper is the enhancement of dissipation by the presence of a nondissipative Hamiltonian term in the vector field. In the context of the Boltzmann equation, the vector field without the Hamiltonian term drives trajectories to the local Maxwell distribution, with the Hamiltonian term to the total Maxwell distribution ([19]). We conjecture that this type of enhancement of dissipation plays an essential role in the emergence of dissipation in dynamics of macroscopic systems. Ignorance of details of a minute irregularity that may arise on microscopic scale is enhanced to the dissipation detectable on the macroscopic scale.
Note also that in the case of infinite-dimensional systems, our calculations are only formal, since we do not specify the function spaces and the smoothness of the functionals. In particular, we do not specify the functional spaces for the state variables, the entropy, the energy, and the dissipation potential. We assume that all functionals are sufficiently smooth (twice Fréchet differentiable) and local (with vanishing boundary conditions).
Among the frameworks, we mention Classical Irreversible Thermodynamics (CIT) [76, 77, 18, 87], Extended Irreversible Thermodynamics [57, 62], Internal Variables Thermodynamics [8], Rational Thermodynamics [94], Rational Extended Thermodynamics [73], Steepest Entropy Ascent [5], and port-Hamiltonian systems [97, 65].
@@
In this paper, we focus on geometric frameworks for multiscale dissipative time evolution. Such frameworks are formulated in the multiscale setting (i.e., they can be used in the context of fluid mechanics, kinetic theory, and other autonomous mesoscopic theories) and provide a unification and additional physical and mathematical insights. Among the geometric frameworks, we discuss generalized gradient dynamics [81, 53], the General Equation for Non-Equilibrium Reversible-Irreversible Coupling (GENERIC) [44, 43, 52, 79, 78, 82, 9], and metriplectic systems [71, 58, 44], where gradient dynamics is combined with Hamiltonian mechanics. In the case when the target mesoscopic theory is not equilibrium thermodynamics but a mesoscopic theory involving fewer details, the role of the entropy is replaced by the Rayleigh dissipation potential [89, 61, 24]. Dissipative terms can be also added as constraints in the variational formulation of mechanics, leading to the d’Alembert variational principle in non-equilibrium thermodynamics [32, 98, 33]. Finally, when we have a mechanical (Hamiltonian) system, dissipative evolution can be constructed from the underlying Poisson bracket and from the Hamiltonian (double bracket dissipation [11] and Ehrenfest regularization [83]).
In particular, we show that the Rayleigh-potential framework can be cast into the GENERIC form, employing generalized gradient dynamics as the dissipative part of the evolution. We also formulate a version of the dissipative d’Alembert variational principle using the Rayleigh dissipation potential, which makes it easier to eliminate fast-relaxing variables in a geometric manner. Finally, we show a variational formulation of the GENERIC framework and a relation of Extended GENERIC with port-Hamiltonian systems [97].
When comparing these frameworks, we focus on the following features: (i) Geometric formulation, since frameworks that are not geometrically consistent may not be invariant with respect to changes of variables; (ii) Modeling minimalism, since approaches that require fewer modeling assumptions are preferable; (iii) Generality, so that framework can be used on various levels of description.
@@
In Section 2, we motivate the frameworks for dissipative evolution by pointing out some problems of CIT. Section 3 contains generalized gradient dynamics while Section 4 is about the Rayleigh dissipation potential. In Section 5 we discuss the dissipative d’Alembert variational principle, and in Section 6 we recall some frameworks where dissipative evolution is constructed from the Poisson bracket. Sections 7 and 8 show the connection of GENERIC with port-Hamiltonian systems and contact geometry, respectively.
@@ 2 Classical irreversible thermodynamics
State variables in the context of Classical Irreversible Thermodynamics [18] are chosen to be $\mathbf{x}=(\mathbf{y},e)$, where $e$ is the volumetric total energy density and $\mathbf{y}$ represents the remaining densities that play the role of state variables (for instance mass density, momentum density, or some internal variable densities). The time evolution of the volumetric energy density $e$ is governed by
@@
where $\mathbf{J}_{e}=e\mathbf{v}+\mathbf{J}_{q}$ is the total energy flux, $\mathbf{v}$ is the barycentric velocity and $\mathbf{J}_{q}$ is the heat flux (see [18] for other definitions of heat flux). This evolution equation expresses the conservation of the total energy.
The evolution equations of the remaining state variables $\mathbf{y}$ are expected in the form
$\partial_{t}\mathbf{y}=-\nabla\cdot(\mathbf{y}\otimes\mathbf{v}+\mathbf{J}_{\mathbf{y}})+\sigma_{\mathbf{y}},$ (2)
where $\mathbf{J}_{\mathbf{y}}$ is the flux of $\mathbf{y}$ and $\sigma_{\mathbf{y}}$ is a source term.
The unknown fluxes and source terms are chosen so that they fulfill the second law of thermodynamics and Onsager-Casimir reciprocal relations. This is achieved by writing the balance equation for the volumetric entropy density $s(\mathbf{x})$ and by rewriting the entropy production as a sum of products of thermodynamic fluxes and forces. The entropy balance equation then reads
$\partial_{t}s(\mathbf{x})=$ $\frac{\partial s}{\partial\mathbf{x}}\cdot\partial_{t}\mathbf{x}=\frac{\partial s}{\partial e}\left(-\nabla\cdot\mathbf{J}_{e}\right)+\frac{\partial s}{\partial\mathbf{y}}\cdot\partial_{t}\mathbf{y}$
@@
$=$ $-\nabla\cdot\left(\mathbf{v}e\frac{\partial s}{\partial e}+\mathbf{J}_{q}\frac{\partial s}{\partial e}+\mathbf{v}\left(\mathbf{y}\cdot\frac{\partial s}{\partial\mathbf{y}}\right)+\mathbf{J}_{\mathbf{y}}\frac{\partial s}{\partial\mathbf{y}}\right)+\mathbf{J}_{e}\cdot\nabla\frac{\partial s}{\partial e}+(\mathbf{v}\cdot\nabla)\frac{\partial s}{\partial\mathbf{y}}\cdot\mathbf{y}+\mathbf{J}_{\mathbf{y}}\cdot\nabla\frac{\partial s}{\partial\mathbf{y}}+\sigma_{\mathbf{y}}\cdot\frac{\partial s}{\partial\mathbf{y}}$
@@
$=$ $-\nabla\cdot\left(\frac{\partial s}{\partial e}\mathbf{J}^{\prime}_{q}+\left(e\frac{\partial s}{\partial e}+\mathbf{y}\cdot\frac{\partial s}{\partial\mathbf{y}}\right)\mathbf{v}\right)+\left(e\nabla\frac{\partial s}{\partial e}+\mathbf{y}\cdot\nabla\frac{\partial s}{\partial\mathbf{y}}\right)\mathbf{v}+\mathbf{J}^{\prime}_{q}\cdot\nabla\frac{\partial s}{\partial e}+\mathbf{J}_{\mathbf{y}}\cdot\nabla\frac{\partial s}{\partial\mathbf{y}}+\sigma_{\mathbf{y}}\cdot\frac{\partial s}{\partial\mathbf{y}}$ (3)
where the derivative of entropy with respect to energy is the inverse temperature, $T=\left(\frac{\partial s}{\partial e}\right)^{-1}$ is the temperature, and the reduced heat flux is defined as $\mathbf{J}^{\prime}_{q}=\mathbf{J}_{q}-\mathbf{J}_{\mathbf{y}}\cdot\left(\frac{\partial e}{\partial\mathbf{y}}\right)_{s}$.
At this point, we need to use the Gibbs-Duhem relation (or extensivity of entropy in terms of the densities), which states that
$e\nabla\frac{\partial s}{\partial e}+\mathbf{y}\cdot\nabla\frac{\partial s}{\partial\mathbf{y}}=0\quad\text{and}\quad e\frac{\partial s}{\partial e}+\mathbf{y}\cdot\frac{\partial s}{\partial\mathbf{y}}=s.$ (4)
Finally, when the entropy derivatives are expressed in terms of the corresponding energy derivatives, the entropy balance equation becomes
@@
$\partial_{t}s=-\nabla\cdot\left(T^{-1}\mathbf{J}^{\prime}_{q}+s\mathbf{v}\right)+\underbrace{\frac{1}{T}\mathbf{J}^{\prime}_{q}\cdot\left(-T^{-1}\nabla\frac{\partial e}{\partial s}\right)+\mathbf{J}_{\mathbf{y}}\cdot\left(-\nabla\frac{\partial e}{\partial\mathbf{y}}\right)+\sigma_{\mathbf{y}}\cdot\left(-\frac{\partial e}{\partial\mathbf{y}}\right)}_{=\sigma_{s}},$ (5)
where the entropy production $\sigma_{s}$ is written as a sum of products of thermodynamic fluxes and forces. The total entropy flux can be identified with the terms under the divergence as $\mathbf{J}_{s}=T^{-1}\mathbf{J}^{\prime}_{q}+s\mathbf{v}$.
@@
To guarantee the non-negativity of the entropy production, the fluxes and source terms are then expressed as linear combinations of the thermodynamic forces (gradients of the derivatives of energy),
@@
$\begin{pmatrix}\mathbf{J}^{\prime}_{q}\\ \mathbf{J}_{\mathbf{y}}\\ \sigma_{s}\end{pmatrix}=\mathbf{M}(\mathbf{x})\cdot\begin{pmatrix}-T^{-1}\nabla\frac{\partial e}{\partial s}\\ -\nabla\frac{\partial e}{\partial\mathbf{y}}\\ -\frac{\partial e}{\partial\mathbf{y}}\end{pmatrix}$ (6)
@@
where $\mathbf{M}$ is a linear operator that is positive semi-definite, $\mathbf{x}\cdot\mathbf{M}(\mathbf{x})\cdot\mathbf{x}\geq 0$ and should satisfy the Onsager-Casimir reciprocal relations.
@@
Entropy is often quadratic in the non-conserved state variables (for instance heat flux in the Cattaneo equation [16]), which leads to Extended Irreversible Thermodynamics [57] and Internal Variable Thermodynamics [8, 7, 68].
@@
  • Evolution of state variables $\mathbf{y}$ does not in general have to be in the divergence form, as in the case of polymeric fluids (conformation tensor is not conserved) [10], magnetohydrodynamics (where the Godunov-Powell terms are important for numerical stability [37]), Reynolds stress in turbulence [51], and others [82]. Instead, the evolution equation for $\mathbf{y}$ can have a part that is in the divergence form and a part that is not. Moreover, we can move some terms between the two parts, which makes the identification of fluxes and forces ambiguous.
@@ 3.1 Generalized gradient dynamics
where $\Xi$ is a dissipation potential, $S$ is the entropy, and $\mathbf{x}^{*}$ are thermodynamic forces conjugate to $\mathbf{x}$. The derivative $\delta\bullet/\delta\bullet$ stands for the functional derivative so that we can differentiate functionals with respect to fields. The dissipation potential is a non-negative functional of $\mathbf{x}^{*}$ that has a minimum at $\mathbf{x}^{*}=0$, is convex in the vicinity of the minimum, and is radially monotonous, $\mathbf{x}^{*}\cdot\frac{\delta\Xi}{\delta\mathbf{x}^{*}}\geq 0$; see [91, 56]. The dissipation potential $\Xi(\mathbf{x},\mathbf{x}^{*})$ also depends on the state variables $\mathbf{x}$.
@@
3.2 Gradient dynamics with constraints
In non-equilibrium thermodynamics, the evolution is typically constrained by energy or momentum conservation. Such constraints can be incorporated into gradient dynamics (7) by means of Lagrange multipliers, which can be geometrically seen as a Morse family [26].
Alternatively, the constraints $c^{\alpha}(\mathbf{x})$ can be incorporated into the framework by forming thermodynamic forces that are orthogonal to the gradients of the constraints [54, 49],
$X^{\alpha}=\mathcal{K}_{\alpha}\mathbf{x}^{*}$ (12)
where $\mathcal{K}$ is a linear operator satisfying $\mathcal{K}_{\alpha}(dc^{\beta})=0\quad\forall\alpha\forall\beta$. When the dissipation potential depends on the conjugate variables only via the thermodynamic forces, $\Xi(\mathbf{x},\mathbf{x}^{*})=\Xi(\mathbf{x},\mathcal{K}\mathbf{x}^{*})$, the evolution equations become
$\dot{\mathbf{x}}=\sum_{\alpha}\mathcal{K}^{\dagger}_{\alpha}\frac{\delta\Xi}{\delta X^{\alpha}}\Big|_{X^{\alpha}=\mathcal{K}_{\alpha}\frac{\delta S}{\delta\mathbf{x}}},$ (13)
where $\mathcal{K}^{\dagger}_{\alpha}$ is the adjoint operator to $\mathcal{K}_{\alpha}$ and where $\frac{\delta\Xi}{\delta X^{\alpha}}\Big|_{X^{\alpha}=\mathcal{K}_{\alpha}\frac{\delta S}{\delta\mathbf{x}}}=J^{\alpha}$ is the thermodynamic flux corresponding to the force $X^{\alpha}$.
Evolution of constraint $c^{\alpha}(\mathbf{x})$ is then given by
$\dot{c}^{\alpha}=\left\langle dc^{\alpha},\dot{\mathbf{x}}\right\rangle=\left\langle dc^{\alpha},\sum_{\beta}\mathcal{K}^{\dagger}_{\beta}\frac{\delta\Xi}{\delta X^{\beta}}\Big|_{X^{\beta}=\mathcal{K}_{\beta}\frac{\delta S}{\delta\mathbf{x}}}\right\rangle=\sum_{\beta}\left\langle\mathcal{K}_{\beta}dc^{\alpha},\frac{\delta\Xi}{\delta X^{\beta}}\Big|_{X^{\beta}=\mathcal{K}_{\beta}\frac{\delta S}{\delta\mathbf{x}}}\right\rangle=0,$ (14)
since $\mathcal{K}_{\beta}dc^{\alpha}=0$ by construction.
For instance, in the case of energy conservation, $\mathbf{x}=e(\mathbf{r})$, we have $\mathcal{K}dE=0$ with $\mathcal{K}=\nabla$ and $dE=1$. In the case of chemical reactions, $\mathbf{x}=(n_{1},n_{2},\dots,n_{N})$ are the number densities of $N$ chemical species and the dynamics is constrained by the admissible chemical reactions defined by the stoichiometric matrix $\nu^{\alpha}_{i}$, where $\alpha$ runs over all reactions and $i$ runs over all species. The constraints represent the conservation of atoms in the chemical reactions and have the form $C=\sum_{i}c_{i}n_{i}$. The operator $\mathcal{K}$ is then given by $\mathcal{K}_{\alpha}\mathbf{x}^{*}=\sum_{i}\nu^{\alpha}_{i}x^{*}_{i}$, from which it follows that $\mathcal{K}_{\alpha}dC=\sum_{i}\nu^{\alpha}_{i}c_{i}=0$ since the number of atoms is conserved in each reaction.